# Towards Hierarchical Rectified Flow

### v1

@@ Abstract
We formulate a hierarchical rectified flow to model data distributions. It hierarchically couples multiple ordinary differential equations (ODEs) and defines a time-differentiable stochastic process that generates a data distribution from a known source distribution. Each ODE resembles the ODE that is solved in a classic rectified flow, but differs in its domain, i.e., location, velocity, acceleration, etc. Unlike the classic rectified flow formulation, which formulates a single ODE in the location domain and only captures the expected velocity field (sufficient to capture a multi-modal data distribution), the hierarchical rectified flow formulation models the multi-modal random velocity field, acceleration field, etc., in their entirety. This more faithful modeling of the random velocity field enables integration paths to intersect when the underlying ODE is solved during data generation. Intersecting paths in turn lead to integration trajectories that are more straight than those obtained in the classic rectified flow formulation, where integration paths cannot intersect. This leads to modeling of data distributions with fewer neural function evaluations. We empirically verify this on synthetic 1D and 2D data as well as MNIST, CIFAR-10, and ImageNet-32 data. Code is available at: https://riccizz.github.io/HRF/.
@@ 2 Preliminaries
Given a dataset ${\cal D}=\{(x_{1})\}$ consisting of samples $x_{1}\sim\rho_{1}$, e.g., images, drawn from an unknown target data distribution $\rho_{1}$, the goal of generative modeling is to learn a model that faithfully captures the dataset distribution $\rho_{1}$ and permits to sample from the learned distribution.
Since we focus primarily on rectified flow, we provide the necessary background in the following. At inference time, rectified flow starts from samples $x_{0}\sim\rho_{0}$ drawn from a known source distribution $\rho_{0}$, e.g., a standard Gaussian. The source distribution samples are pushed forward from time $t=0$ to target time $t=1$ via integration along a trajectory specified via a learned velocity field $v(z_{t},t)$. This learned velocity field depends on the current time $t$ and the sample location $z_{t}$ at time $t$. Formally, we obtain samples by numerically solving the ordinary differential equation (ODE)
@@
To learn the velocity field, at training time, rectified flow constructs random pairs $(x_{0},x_{1})$, consisting of a source distribution sample $x_{0}\sim\rho_{0}$ and a target distribution sample $x_{1}\sim{\cal D}$ drawn from a given dataset ${\cal D}$ consisting of samples which are assumed to be drawn from the unknown target distribution $\rho_{1}$. For a uniformly drawn time $t\sim U[0,1]$, the time-dependent location $x_{t}$ is computed from the pair $(x_{0},x_{1})$ using linear interpolation of $(x_{0},x_{1})$, i.e.,
@@ 3.1 Velocity distribution and case study with Gaussian Mixtures
The linear interpolation in Eq. 2 defines a time-differentiable stochastic process with the random velocity field $v(x_{t},t)=x_{1}-x_{0}$, where $x_{0}\sim\rho_{0}$ and $x_{1}\sim\rho_{1}$. Note, the source and target distributions are independent. The following theorem characterizes the distribution of the velocity at a specific space time location $x_{t}$,
@@ 3.2 Modeling the Velocity Distribution
To accurately model the “ground-truth” velocity distribution, we can use rectified flow for velocities rather than locations, which are used in the classic rectified flow formulation. This is equivalent to learning the acceleration. To see this, first, consider classic rectified flow again: we construct a time-dependent location $x_{t}$ from pairs $(x_{0},x_{1})$, compute the “ground-truth” velocity $v(x_{t},t)=\partial x_{t}/\partial t$, and train a velocity model $v_{\theta}(x_{t},t)$ to match this “ground-truth” velocity $v(x_{t},t)$.
@@
Using both the source velocity sample $v_{0}$ and the target velocity sample $v_{1}(x_{t},t)$, and following classic rectified flow, we introduce a new time-axis $\tau\in[0,1]$ and construct a time-dependent velocity $v_{\tau}(x_{t},t)=(1-\tau)v_{0}+\tau v_{1}(x_{t},t)$ at time $t$ and at location $x_{t}$. Using it, we obtain the “ground-truth” acceleration from the time-dependent velocity $v_{\tau}(x_{t},t)$ via $a(x_{t},t,v_{\tau},\tau)=\partial v_{\tau}/\partial\tau=v_{1}(x_{t},t)-v_{0}=x_{1}-x_{0}-v_{0}$.
@@
Here, $a(x_{t},t,u_{\tau},\tau)=\mathbb{E}_{\pi_{0},\pi_{1}(v;x_{t},t)}[V_{1}-V_{0}|V_{\tau}=u]=\mathbb{E}_{\pi_{0},\rho_{0},\rho_{1}}[X_{1}-X_{0}-V_{0}|V_{\tau}=u,X_{t}=x_{t}]$ is the expected acceleration vector field.
@@ 4       Perform gradient update on $\theta$
and $g$ is a deterministic function. The randomness comes from the initial conditions for data and velocity. This is different from the sampling process governed by a stochastic differential equation (SDE) used in diffusion models Song et al. (2021b), where the randomness comes from the Wiener process and the initial condition for data.
@@ 5 Related Work
Generative Modeling: GANs (Goodfellow et al., 2014; Arjovsky et al., 2017), VAEs (Kingma & Welling, 2014), and normalizing flows (Tabak & Turner, 2013; Rezende & Mohamed, 2015; Dinh et al., 2017; Huang et al., 2018; Durkan et al., 2019) are classic methods for learning deep generative models. GANs excel in generating high-quality images but face challenges like training instability and mode collapse due to their min-max update mechanism. VAEs and normalizing flows rely on maximum likelihood estimation (MLE) for training, which necessitates architectural constraints or special approximations to ensure manageable likelihood computations. VAEs often employ a conditional Gaussian distribution alongside variational approximations, while the discrete normalizing flows utilize specifically designed invertible architectures and require costly Jacobian matrix calculations. Extending the discrete normalizing flow to continuous cases enabled the Jacobian to be unstructured yet estimable using trace estimation methods (Hutchinson, 1990; Chen et al., 2018; Grathwohl et al., 2019). However, using maximum likelihood estimation (MLE) for this mapping requires costly backpropagation through numerical integration. Regulating the path can minimize solver calls (Finlay et al., 2020; Onken et al., 2021), but it doesn’t resolve the fundamental optimization challenges. Rozen et al. (2021); Ben-Hamu et al. (2022) considered simulation-free training by fitting a velocity field, but still present scalability issues (Rozen et al., 2021) and biased optimization (Ben-Hamu et al., 2022).
Recent research has utilized diffusion processes, particularly the Ornstein-Uhlenbeck (OU) process, to link the target distribution $\rho_{1}$ with a source distribution $\rho_{0}$. This involves a stochastic differential equation (SDE) that evolves over infinite time, framing generative model learning as fitting the reverse evolution of the SDE from Gaussian noise to $\rho_{1}$ (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b). This method learns the velocity field by estimating the score function $\nabla\log(\rho_{t}(x))$ using Fischer divergence instead of the maximum likelihood estimation (MLE) objective. Although the diffusion models have demonstrated significant potential for modeling high-dimensional distributions (Rombach et al., 2022; Hoogeboom et al., 2022; Saharia et al., 2022), the requirement for infinite time evolution, heuristic time step parameterization (Xiao et al., 2022), and the unclear significance of noise and score (Bansal et al., 2024; Lu et al., 2022) pose challenges. Notably, the score-based diffusion models typically require a large number of time steps to generate data samples. In addition, calculating the actual likelihoods necessitates using the ODE probability flow linked to the SDE (Song et al., 2021b). These highlight the need for further exploration of effective ODE-driven methods for learning the data distribution.
@@ 6 Discussion & Conclusion
Acknowledgements: Work supported in part by NSF grants 1934757, 2008387, 2045586, 2106825, MRI 1725729, NIFA award 2020-67021-32799, and an Alfred P. Sloan research fellowship.
@@ Appendix B Proof of Corollary 1
Bromiley (2003) summarizes a few useful properties of the product and convolution of the Gaussian distributions. We state the relevant results here for our proof of Corollary 1.
@@
Proof of Corollary 1: We first compute the density of $X_{t}$ using Theorem 1 using the specific $\rho_{0}$ and $\rho_{1}$.
@@ Appendix C Proof of Theorem 2
According to Theorem 3.3 of Liu et al. (2023), the ODE in Eq. 7 generates the samples from the ground-truth velocity distributions at space time location $(x_{t},t)$.
We consider the characteristic function of $Z_{t+\Delta t}=Z_{t}+V\Delta t$ for $t\in[0,1]$ and $\Delta t\in[0,1-t]$, assuming that $Z_{t}$ has the same distribution as $X_{t}$. If the characteristic functions of $Z_{t+\Delta t}$ and $X_{t+\Delta t}$ agree, then we can show that $Z_{t+\Delta t}$ and $X_{t+\Delta t}$ enjoy the same distributions. In addition, since $Z_{0}$ and $X_{0}$ have the same distributions, we can conclude that $Z_{t}$ and $X_{t}$ have the same distributions for $t\in[0,1]$.
@@ Appendix D Density estimation
Algorithm 4 outlines the procedure for the likelihood computation with a randomly drawn $t\in(0,1)$. Optionally, we can average across randomly drawn $t\in(0,1)$.
@@ 3 Compute $z_{t}=tz_{1}+(1-t)z_{0}$ and $u=\frac{z_{1}-z_{t}}{1-t}$ ;
In Table 1, we report bits per dimension (bpd) for experiments on the 1D $1\mathcal{N}\to 2\mathcal{N}$, 2D $8\mathcal{N}\to$ moon, CIFAR-10, and ImageNet-32 data. For 1D data, $z_{0}=0$ suffices for compelling results. For higher dimensional data, we use $N=20$ $z_{0}$ as shown in the optional line 4 of Algorithm 3 to compute the bits per dimension. We observe that HRF2 consistently outperforms the RF baseline.
@@ F.1 Low Dimensional Experiments
For 1D and 2D experiments, we use the same neural network. It consists of two parts. The first part processes the space and time information separately using a Sinusoidal Positional Embedding and linear layers. In the second part, the processed information is concatenated and passed through a series of linear layers to produce the final output. Compared to the baseline, our HRF model with depth $D$ takes $D$ times more space and time information as input. Therefore, the first part of the network has $D$ times more embedding and linear layers to handle the spatial and temporal information from different depths. However, by adjusting the dimensions of the hidden layers, the size of our network is only one-fourth that of the baseline, while achieving superior performance. For each dataset in the low-dimensional experiments, we use 100,000 data points for training and another 100,000 data points for evaluation. For each set of experiments, we train five different models using five random seeds. During evaluation, we conduct a total of 125 experiments and average the results to ensure the fairness and validity of our findings.
@@ F.2 High Dimensional Experiments
MNIST. For MNIST, we use a single U-Net and modify the ResNet block. Similar to the neural network used in our low-dimensional experiments, each ResNet block has two parts. In the first part, we handle 2 sets of space-time information, i.e., $(x_{t},t)$ and $(v_{\tau},\tau)$, separately with 2 distinct pathways: convolutional layers for spatial data and linear layers for time embeddings. In the second part, all the spatial data and time embeddings are added together and passed through a series of linear layers to capture the space-time dependencies. For a fair evaluation, we adjusted the number of channels such that the model sizes approximately match (ours: 1.07M parameters vs. baseline: 1.08M parameters). We note that the HRF formulation significantly outperforms the baseline. Results were shown in Fig. 6. More results are provided in Section H.3.

### v2

@@ Abstract
We formulate a hierarchical rectified flow to model data distributions. It hierarchically couples multiple ordinary differential equations (ODEs) and defines a time-differentiable stochastic process that generates a data distribution from a known source distribution. Each ODE resembles the ODE that is solved in a classic rectified flow, but differs in its domain, i.e., location, velocity, acceleration, etc. Unlike the classic rectified flow formulation, which formulates a single ODE in the location domain and only captures the expected velocity field (sufficient to capture a multi-modal data distribution), the hierarchical rectified flow formulation models the multi-modal random velocity field, acceleration field, etc., in their entirety. This more faithful modeling of the random velocity field enables integration paths to intersect when the underlying ODE is solved during data generation. Intersecting paths in turn lead to integration trajectories that are more straight than those obtained in the classic rectified flow formulation, where integration paths cannot intersect. This leads to modeling of data distributions with fewer neural function evaluations. We empirically verify this on synthetic 1D and 2D data as well as MNIST, CIFAR-10, and ImageNet-32 data. Our code is available at: https://riccizz.github.io/HRF/.
@@ 2 Preliminaries
Given a dataset ${\cal D}=\{(x_{1})\}$ consisting of samples $x_{1}\sim\rho_{1}$, e.g., images, drawn from an unknown target data distribution $\rho_{1}$, the goal of generative modeling is to learn a model that faithfully captures the unknown target data distribution $\rho_{1}$ and permits to sample from the learned distribution.
Since we focus primarily on rectified flow, we provide its formulation in the following. At inference time, rectified flow starts from samples $x_{0}\sim\rho_{0}$ drawn from a known source distribution $\rho_{0}$, e.g., a standard Gaussian. The source distribution samples are pushed forward from time $t=0$ to target time $t=1$ via integration along a trajectory specified via a learned velocity field $v(z_{t},t)$. This learned velocity field depends on the current time $t$ and the sample location $z_{t}$ at time $t$. Formally, we obtain samples by numerically solving the ordinary differential equation (ODE)
@@
To learn the velocity field, at training time, rectified flow constructs random pairs $(x_{0},x_{1})$, consisting of a source distribution sample $x_{0}\sim\rho_{0}$ and a target distribution sample $x_{1}\sim{\cal D}$. The latter is drawn from a given dataset ${\cal D}$ consisting of samples which are assumed to be drawn from the unknown target distribution $\rho_{1}$. For a uniformly drawn time $t\sim U[0,1]$, the time-dependent location $x_{t}$ is computed from the pair $(x_{0},x_{1})$ using linear interpolation of $(x_{0},x_{1})$, i.e.,
@@ 3.1 Velocity distribution and case study with Gaussian Mixtures
The linear interpolation in Eq. 2 defines a time-differentiable stochastic process with the random velocity field $v(x_{t},t)=x_{1}-x_{0}$, where $x_{0}\sim\rho_{0}$ and $x_{1}\sim\rho_{1}$. Note, the source and target distributions are independent. The following theorem characterizes the distribution of the velocity at a specific space time location $(x_{t},t)$:
@@ 3.2 Modeling the Velocity Distribution
To accurately model the “ground-truth” velocity distribution, we can use rectified flow for velocities rather than locations, which are used in the classic rectified flow formulation. This is equivalent to learning the acceleration. To see this, first, consider classic rectified flow again: we construct a time-dependent location $x_{t}$ from pairs $(x_{0},x_{1})$, compute the “ground-truth” velocity $v_{\mathrm{gt}}(x_{t},t)=\partial x_{t}/\partial t$, and train a velocity model $v_{\theta}(x_{t},t)$ to match this “ground-truth” velocity $v_{\mathrm{gt}}(x_{t},t)$.
@@
Using both the source velocity sample $v_{0}$ and the target velocity sample $v_{1}(x_{t},t)$, and following classic rectified flow, we introduce a new time-axis $\tau\in[0,1]$ and construct a time-dependent velocity $v_{\tau}(x_{t},t)=(1-\tau)v_{0}+\tau v_{1}(x_{t},t)$ at time $t$ and at location $x_{t}$. Using it, we obtain the “ground-truth” acceleration from the time-dependent velocity $v_{\tau}(x_{t},t)$ via $a_{\mathrm{gt}}(x_{t},t,v_{\tau},\tau)=\partial v_{\tau}/\partial\tau=v_{1}(x_{t},t)-v_{0}=x_{1}-x_{0}-v_{0}$.
@@
Here, $a(x_{t},t,u_{\tau},\tau)\!=\!\mathbb{E}_{\pi_{0},\pi_{1}(v;x_{t},t)}[V_{1}\!-\!V_{0}|V_{\tau}\!=\!u_{\tau}]\!=\!\mathbb{E}_{\pi_{0},\rho_{0},\rho_{1}}[X_{1}\!-\!X_{0}\!-\!V_{0}|V_{\tau}\!=\!u_{\tau},X_{t}\!=\!x_{t}]$ is the expected acceleration vector field.
@@ 4       Perform gradient update on $\theta$
and $g$ is a deterministic function. The randomness comes from the initial conditions for data and velocity. This is different from the sampling process governed by a stochastic differential equation (SDE) used in diffusion models (Song et al., 2021b), where the randomness comes from the Wiener process and the initial condition for data.
@@ 5 Related Work
Generative Modeling: GANs (Goodfellow et al., 2014; Arjovsky et al., 2017), VAEs (Kingma & Welling, 2014), and normalizing flows (Tabak & Turner, 2013; Rezende & Mohamed, 2015; Dinh et al., 2017; Huang et al., 2018; Durkan et al., 2019) are classic methods for learning deep generative models. GANs excel in generating high-quality images but face challenges like training instability and mode collapse due to their min-max update mechanism. VAEs and normalizing flows rely on maximum likelihood estimation (MLE) for training, which necessitates architectural constraints or special approximations to ensure manageable likelihood computations. VAEs often employ a conditional Gaussian distribution alongside variational approximations, while the discrete normalizing flows utilize specifically designed invertible architectures and require costly Jacobian matrix calculations. Extending the discrete normalizing flow to continuous cases enabled the Jacobian to be unstructured yet estimatable using trace estimation methods (Hutchinson, 1990; Chen et al., 2018; Grathwohl et al., 2019). However, using maximum likelihood estimation (MLE) for this mapping requires costly backpropagation through numerical integration. Regularizing the path can minimize solver calls (Finlay et al., 2020; Onken et al., 2021), but it doesn’t resolve the fundamental optimization challenges. Rozen et al. (2021); Ben-Hamu et al. (2022) considered simulation-free training by fitting a velocity field, but still present scalability issues (Rozen et al., 2021) and biased optimization (Ben-Hamu et al., 2022).
Recent research has utilized diffusion processes, particularly the Ornstein-Uhlenbeck (OU) process, to link the target distribution $\rho_{1}$ with a source distribution $\rho_{0}$. This involves a stochastic differential equation (SDE) that evolves over infinite time, framing generative model learning as fitting the reverse evolution of the SDE from Gaussian noise to $\rho_{1}$ (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b). This method learns the velocity field by estimating the score function $\nabla\log(\rho_{t}(x))$ using the Fischer divergence instead of the maximum likelihood estimation (MLE) objective. Although diffusion models have demonstrated significant potential for modeling high-dimensional distributions (Rombach et al., 2022; Hoogeboom et al., 2022; Saharia et al., 2022), the requirement for infinite time evolution, heuristic time step parameterization (Xiao et al., 2022), and the unclear significance of noise and score (Bansal et al., 2024; Lu et al., 2022) pose challenges. Notably, the score-based diffusion models typically require a large number of time steps to generate data samples. In addition, calculating the actual likelihoods necessitates using the ODE probability flow linked to the SDE (Song et al., 2021b). These highlight the need for further exploration of effective ODE-driven methods for learning the data distribution.
@@ 6 Discussion & Conclusion
Acknowledgements: Work supported in part by NSF grants 1934757, 2008387, 2045586, 2106825, MRI 1725729, NIFA award 2020-67021-32799, and the Alfred P. Sloan Foundation.
@@ Appendix B Proof of Corollary 1
Bromiley (2003) summarizes a few useful properties for the product and convolution of Gaussian distributions. We state the relevant results here for our proof of Corollary 1.
@@
Proof of Corollary 1: We first compute the density of $X_{t}$ using Theorem 1 with the specific $\rho_{0}$ and $\rho_{1}$:
@@ Appendix C Proof of Theorem 2
According to Theorem 3.3 of Liu et al. (2023), the ODE in Eq. 7 generates the samples from the ground-truth velocity distributions at space time location $(x_{t},t)$. In other words, the random variable $V\sim\pi_{1}$.
Now we consider the characteristic function of $Z_{t+\Delta t}=Z_{t}+V\Delta t$ for $t\in[0,1]$ and $\Delta t\in[0,1-t]$, assuming that $Z_{t}$ has the same distribution as $X_{t}$. If the characteristic functions of $Z_{t+\Delta t}$ and $X_{t+\Delta t}$ agree, then $Z_{t+\Delta t}$ and $X_{t+\Delta t}$ have the same distribution.
@@ Appendix D Density estimation
where $z^{(i)}_{0}\sim\rho_{0}$. Algorithm 4 outlines the procedure for the likelihood computation with a randomly drawn $t\in(0,1)$. Optionally, we can average across randomly drawn $t\in(0,1)$.
@@ 3 Compute $z_{t}=tz_{1}+(1-t)z_{0}$ and $u=\frac{z_{1}-z_{t}}{1-t}$ ;
In Table 1, we report bits per dimension (bpd) for experiments on the 1D $1\mathcal{N}\to 2\mathcal{N}$, 2D $8\mathcal{N}\to$ moon, CIFAR-10, and ImageNet-32 data. For 1D data, $z_{0}=0$ suffices for compelling results. For higher dimensional data, we use $20$ samples of $z_{0}$ as shown in the optional line 4 of Algorithm 3 to compute the bits per dimension. We observe that HRF2 consistently outperforms the RF baseline.
@@ F.1 Low Dimensional Experiments
For the 1D and 2D experiments, we use the same neural network. It consists of two parts. The first part processes the space and time information separately using sinusoidal positional embedding and linear layers. In the second part, the processed information is concatenated and passed through a series of linear layers to produce the final output. Compared to the baseline, our HRF model with depth $D$ takes $D$ times more space and time information as input. Therefore, the first part of the network has $D$ times more embedding and linear layers to handle spatial and temporal information from different depths. However, by adjusting the dimensions of the hidden layers, we reduced the network size to just one-fourth of the baseline, while achieving superior performance. For each dataset in the low-dimensional experiments, we use 100,000 data points for training and another 100,000 data points for evaluation. For each set of experiments, we train five different models using five random seeds. During the evaluation, we performed a total of 125 experiments and averaged the results to ensure the fairness and validity of our findings.
@@ F.2 High Dimensional Experiments
MNIST. For MNIST, we use a single U-Net and modify the ResNet block. Similar to the neural network used in our low-dimensional experiments, each ResNet block has two parts. In the first part, we handle two sets of space-time information, i.e., $(x_{t},t)$ and $(v_{\tau},\tau)$, separately with 2 distinct pathways: convolutional layers for spatial data and linear layers for time embeddings. In the second part, all the spatial data and time embeddings are added together and passed through a series of linear layers to capture the space-time dependencies. For a fair evaluation, we adjusted the number of channels such that the model sizes approximately match (ours: 1.07M parameters vs. baseline: 1.08M parameters). We note that the HRF formulation significantly outperforms the baseline. The results were shown in Fig. 6. More results are provided in Section H.3.